# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ VIII-E Network Intelligence Technologies for 6G
VIII-E1 Intelligent Core
  • Vision: In the core network, 6G will go beyond 5G’s softwarization by adopting intelligence via pervasive AI/ML. This can be seen as an evolution from cloud to cloud intelligence.
  • Introduction: Mobile networks consist of three main components: core network, RAN, and mobile devices, as illustrated in Figure 27. The core and RAN operate cooperatively between the Internet and mobile devices. The core network is responsible for the overall network operation, management, and security, whereas the RAN handles wireless access to serve mobile devices. For example, one of the key tasks of the core network is to guarantee secure and reliable end-to-end communication between source and destination nodes. Today’s core networks are software-based and cloud-native. The next major step in the evolution of core networks is the adoption of intelligence via AI/ML in the 6G era. The intelligent network core is based on the pervasive utilization of AI/ML in order to improve the design, operation, management, maintenance, and security of the network. DL is an attractive technology for the core since it achieves excellent performance when the training datasets are large. Data acquisition is a crucial component of the intelligent core.
  • Past and Present: Each mobile network generation has its own core network that responds to the demands and expectations of that particular generation. Since the requirements of mobile networks are constantly growing, the capabilities of the core networks evolve generation by generation. In short, the core networks have evolved from 2G’s circuit-switched network through 3G’s and 4G’s packet cores to the latest software/service-based architecture of 5G. Specifically, a new cloud-native service-oriented core network was developed for 5G, as defined by 3GPP in Release 15 [1]. The main features of the 5G Core include SBA, SDN, NFV, and network slicing. SBA relies on a vast variety of interconnected network functions, each of which provides one or more services that are accessible by other network functions. This architecture allows the agile addition of novel functionalities. SDN enables cloud-native and flexible network management. NFV and network slicing enable the division of the network resources into independent virtual slices, each of which can be dedicated to a specific service. Virtual slicing facilitates the introduction of new services and tailored customer solutions.
  • Opportunities and Challenges: While softwarization made 5G’s core network cloud native, AI/ML is expected to make 6G’s core intelligent and take mobile networks to the era of intelligence. AI/ML can potentially improve the operation, management, and security of the network. Although much research has been conducted, there are many issues to be resolved in harnessing the synergy of AI/ML and mobile networks. In [179, 460, 178, 461, 462, 463], many challenges and future directions were discussed. These can be divided into several categories: infrastructure, AI/ML algorithms, network optimization, core-edge cooperation, and security. First of all, network infrastructure needs to be highly capable from the core to the edge (in terms of communication, computation, and information exchange) to enable the pervasive and efficient use of AI/ML throughout the network. There is a challenging trade-off between the cost and performance in the update of the infrastructure.
@@ VIII-E1 Intelligent Core
  • Literature and Future Directions: In the literature, mobile network intelligence has been extensively studied since the late 2010s, with a focus on 6G networks. Recent research progress and open issues regarding intelligent networking have been reviewed in numerous survey articles [179, 464, 460, 178, 465, 461, 462, 466, 181, 467, 468, 469, 463, 470, 471]. In [179, 178, 465, 181], the works focused on DL, discussing its applications in wireless networks to achieve network intelligence. The paper [179] surveyed DL applications for PHY, data link, and routing layers. In [178], a thorough overview was provided of DL in mobile and wireless networking, with the main focus on the technological enablers of DL in networking, DL-driven network aspects, and customizing DL for networks.
@@
    The tutorial in [468] presented an architectural framework for 6G holistic network virtualization and pervasive network intelligence, reviewing also topics like the interplay between network slicing and digital twins, AI for networking, and networking for AI. In [469], the work introduced a concept of intelligence-endogenous networks for 6G, utilizing AI and knowledge graph technologies. The article [463] surveyed ML-based resource management for 6G networks. The covered aspects include resource allocation, task offloading, mobility management, energy efficiency maximization, and latency minimization. In [470], the authors provided a comprehensive survey on AI for 6G networks, focusing on AI-aided technologies and applications. The work [471] discussed generative AI for 6G wireless intelligence. Future research directions were discussed earlier in the opportunities and challenges part of this section.
Figure 28: Six levels of edge intelligence [472].
VIII-E2 Intelligent Edge
@@
  • Introduction: Intelligent edge, commonly known as edge intelligence (EI), refers to a combination of edge computing and AI, benefiting from the synergy of both technologies [472]. Edge computing drives computational capabilities, tasks, and applications from the core to the network edge [472]. The network edge consists of different types of edge nodes, typical ones being edge servers and end-devices (e.g., mobile and IoT devices) [472]. The edge computing paradigm shift, which moves computation resources closer to the edge information sources, naturally leads to reduced latency and communication overhead.
@@
  • Past and Present: While EI is a new concept, its components, AI and edge computing, have been studied separately in the past. AI research started already in the 1950s, with ups and downs on the road, until it reached a level of constant progress through the introduction of DL in the 2000s. Due to advances in other key technologies, AI boomed in the 2010s, reaching also mobile networks. AI was introduced for 5G in 3GPP’s Release 18 [18]. The history of AI was discussed in more detail in Section VII. Although it took decades for AI to mature, edge computing has evolved rapidly. Specifically, edge computing became a popular topic in the 2010s, attracting wide interest in academia and industry. Consequently, edge computing has been supported by 5G since Release 15 [1], with constant enhancements in Release 16 [13], 17 [17], and 18 [18]. 3GPP is continuing its work on edge computing toward EI. The term EI was first introduced in the white paper of the Gartner Hype Cycle in 2018 [472]. Since then, EI has been one of the key 6G topics in academia and industry, spurring a broad range of publications, projects, collaborations, and development efforts.
@@
  • Opportunities and Challenges: Combining edge computing and AI/ML is a natural step in the evolution of mobile networks toward pervasive network intelligence. Edge computing offers a computation- and data-intensive platform, with the beneficial features of low latency, reduced communication overhead, high energy efficiency, protected privacy, and context awareness. AI has the ability to free the potential offered by edge computing by optimizing the network edge (e.g., performance, efficiency, adaptability, automation, and security/privacy) and enabling a vast range of disruptive applications. Due to its versatile benefits, EI is expected to play a major role in the 6G network intelligence.
@@
    To realize large-scale EI in practice, fundamental challenges must be addressed and resolved. At a high level, the main challenges are related to the computation infrastructure, AI/ML models/algorithms, data availability, network management, network security, and privacy protection. First, large-scale EI requires a ubiquitous computation infrastructure, with powerful computation and storage capabilities at the edge servers, which is a highly non-trivial trade-off between cost and performance. For each EI purpose/task, an efficient AI/ML approach, with practical data acquisition mechanisms, must be developed. Since EI will change the threat landscape, novel security approaches are required. Due to massive amounts of generated data, EI will be vulnerable to privacy threats, calling for advanced privacy protection techniques. For the success of EI, all the aforementioned issues must be properly addressed. Comprehensive discussions of open challenges and future directions can be found in [472, 473, 474, 475, 182].
@@
VIII-E3 Intelligent Air Interface
@@
    Figure 29: Three evolution phases toward AI-native air interface [487].
@@
    TABLE IX: Summary of beyond-communication technologies for 6G
@@
  • Literature and Future Directions: In the literature, AI/ML-enhanced air interface has been studied since the mid-2010s. The main focus has been on DL-aided PHY layer designs, ranging from enhancing/replacing single or multiple processing blocks (such as channel coding, symbol mapping, channel estimation, equalization, detection, decoding, and symbol demapping) to replacing the entire PHY layer with an end-to-end learning process [487]. Furthermore, AI/ML-assisted MAC layer design has been widely examined, often related to resource allocation problems with DRL approaches [488]. Due to the immaturity of this field of research, major efforts are needed in all of these directions in the future, including AI/ML assistance from smaller to larger entities.
@@
    The AI/ML-assisted air interface has been reviewed in the recent literature [490, 491, 492, 493, 494, 495, 496, 497, 489, 487, 183, 498, 186, 499, 500]. In general, these surveys cover the main literature, recent advances, and open challenges. Most studies considered PHY layer design using DL [490, 491, 492, 493, 494, 495, 496, 497, 183, 186, 499, 500]. These DL works can be further divided into the 6G-specific [183, 186, 499], B5G/5G-specific [495, 497], and generic surveys [490, 491, 492, 493, 494, 496, 500]. Additionally, federated learning was reviewed for the PHY layer in [498]. The design of the MAC layer functions was discussed in [494, 489, 487]. The work [487] provided a vision toward an AI/ML-native 6G air interface, ideally enabling optimized communication schemes for any hardware, wireless environment, and application. The paper described the main transition phases on the road to achieving AI/ML-nativeness and discussed the needed PHY/MAC layer learning procedures. In [489], the authors introduced a framework for an AI/ML-enabled 6G air interface, covering PHY/MAC layer designs and interactions with the higher layers. Major research efforts are required in the near future to realize the vision of an intelligent 6G air interface. The corresponding challenges were discussed earlier. In particular, the 6G-specific characteristics and requirements must be properly addressed in the design of AI/ML-enhanced air interface.
VIII-F Beyond-Communication Technologies for 6G
@@
The integration of beyond-communication technologies will be a major paradigm shift in the evolution of mobile networks, significantly extending the way how mobile networks can be exploited. The main beyond-communication technologies include computation, sensing, and energy, as shown in Figure 30. Currently, 5G networks have adopted edge computing capabilities and enhanced positioning in its palette of technologies. This paves the way for the broader usage of beyond-communication technologies in the 6G era. Beyond-communication technologies are expected to enable a broad range of novel capabilities and applications for 6G. In the following, we discuss integrated communication, computation, and caching (i3C), ISAC, and WET. These technologies are summarized in Table IX.
@@
VIII-F1 Integrated Communication, Computation, and Caching
@@
  • Introduction: i3C refers to the joint design of 3C technologies at the edge of the network [501]. Through joint design, 3C technologies can be used to benefit each other, enhancing the performance and extending the capabilities of mobile networks. Consequently, i3C enables novel services and applications. i3C is considered a major paradigm shift in the design of mobile networks [502].
@@
  • Literature and Future Directions: In the literature, i3C has been actively studied since the mid-2010s. Recent survey papers have reviewed the latest achievements and related literature [501, 504, 505, 506, 502, 503]. In [501], a survey was provided on mobile edge networks from the perspective of the 3C convergence. Computing and caching at the edge were first reviewed individually, and then the synergy between the 3C technologies was discussed. The work [504] discussed FL-assisted MEC, communication, and caching to provide more intelligence to the network edge. In [505], 3C resource sharing was explored for D2D IoT scenarios. The authors in [506] presented a review of the 3C convergence in IoT. In [502], a comprehensive survey was provided on the integration of 3C and control (i4C) for beyond 5G networks. The study [503] gave a thorough discussion on i3C in terms of cloud-edge cooperation, resource management, and intelligence.
@@
VIII-F2 Integrated Sensing and Communication
@@
    Due to the different nature of sensing and communication, cellular ISAC has certain shortcomings in practice [509]. First, ISAC requires a full-duplex operation or equivalent. Second, the sensing distance may be limited due to the limited transmission power of the cellular BS. Third, performance trade-offs exist between sensing and communication due to their conflicting targets and requirements. Typical trade-offs are defined as the PHY layer, spatial degrees of freedom, and cross-layer trade-offs, which are further discussed in [507].
@@
  • Literature and Future Directions: Since the late 2010s, ISAC has been one of the main 6G topics in the literature. Over the years, a vast range of aspects has been studied, such as information theoretical limits, performance trade-offs, communication-assisted sensing, sensing-assisted communication (beamforming, resource allocation, PHY layer security), signal processing (waveform optimization, MIMO design, receiver processing), network management (resource allocation, higher layer designs), integration with other emerging technologies (AI/ML, THz, RISs, V2X, UAVs, satellites), and potential applications (smart home/factory/city/healthcare, digital twins, XR, vehicular systems, remote sensing, environmental monitoring, etc.).
    The aforementioned topics and the related literature have been thoroughly reviewed in numerous surveys [514, 509, 510, 511, 515, 508, 228, 513, 507, 516, 517, 518, 519, 520, 521, 234, 522, 523, 524, 525, 526, 527, 512, 238, 240, 528, 529, 530, 531, 532, 533, 534, 535]. At a high level, these survey papers can be classified into three categories: generic, specific, and 6G-oriented. In [514, 509, 508, 513, 516, 525, 528], the generic surveys discussed ISAC from a wide perspective, covering topics such as fundamentals, state-of-the-art designs, recent advancements, potential applications, open problems, and future research guidelines. The specific surveys reviewed narrower aspects of ISAC, including fundamental limits [513], channel modeling [529], waveform design [517], ISAC signals [526], THz ISAC [228, 234, 238, 240], ISAC for IoT [515], RIS-assisted ISAC [521, 522, 523], holographic ISAC [520], UAV-assisted ISAC [524], ISAC for vehicular communication networks [518], and AI/ML-assisted ISAC [519, 527]. The 6G-oriented surveys provided comprehensive explorations of ISAC in the context of 6G [510, 511, 507, 512, 530, 531, 532, 533, 534, 535].
@@
VIII-F3 Wireless Energy Transfer

### v3

@@ IX-E1 INTELLIGENT CORE
  • Past and Present: Each mobile network generation has its own core network that responds to the demands and expectations of that particular generation. Since the requirements of mobile networks are constantly growing, the capabilities of the core networks evolve generation by generation. In short, the core networks have evolved from the circuit-switched network of 2G through the packet cores of 3G and 4G to the latest software/service-based architecture of 5G. Specifically, a new cloud-native service-oriented core network was developed for 5G, as defined by 3GPP in Release 15 [1]. The main features of the 5G Core include SBA, SDN, NFV, and network slicing. SBA relies on a vast variety of interconnected network functions, each of which provides one or more services that are accessible by other network functions. This architecture allows the agile addition of novel functionalities. SDN enables cloud-native and flexible network management. NFV and network slicing enable the division of the network resources into independent virtual slices, each of which can be dedicated to a specific service. Virtual slicing facilitates the introduction of new services and tailored customer solutions.
  • Opportunities and Challenges: While softwarization made 5G core network cloud native, AI/ML is expected to make 6G core intelligent and take mobile networks to the era of intelligence. AI/ML can potentially improve the operation, management, and security of the network. Although much research has been conducted, there are many issues to be resolved in harnessing the synergy of AI/ML and mobile networks. In [197, 473, 196, 474, 475, 476], many challenges and future directions were discussed. These can be divided into several categories: infrastructure, AI/ML algorithms, network optimization, core-edge cooperation, and security. First of all, network infrastructure needs to be highly capable from the core to the edge (in terms of communication, computation, and information exchange) to enable the pervasive and efficient use of AI/ML throughout the network. There is a challenging trade-off between the cost and performance in the update of the infrastructure.
@@
  • Literature and Future Directions: In the literature, mobile network intelligence has been extensively studied since the late 2010s, with a focus on 6G networks. Recent research progress and open issues regarding intelligent networking have been reviewed in numerous survey articles [477, 473, 196, 478, 474, 475, 479, 199, 480, 481, 482, 476, 483, 484]. In [196, 478, 199], the works focused on DL, discussing its applications in wireless networks to achieve network intelligence. In [196], a thorough overview was provided of DL in mobile and wireless networking, with the main focus on the technological enablers of DL in networking, DL-driven network aspects, and customizing DL for networks.
@@
    The tutorial in [481] presented an architectural framework for 6G holistic network virtualization and pervasive network intelligence, reviewing also topics like the interplay between network slicing and digital twins, AI for networking, and networking for AI. In [482], the work introduced a concept of intelligence-endogenous networks for 6G, utilizing AI and knowledge graph technologies. The article [476] surveyed ML-based resource management for 6G networks. The covered aspects included resource allocation, task offloading, mobility management, energy efficiency maximization, and latency minimization. In [483], the authors provided a comprehensive survey on AI for 6G networks, focusing on AI-aided technologies and applications. The work [484] discussed generative AI for 6G wireless intelligence. Future research directions were discussed earlier in the opportunities and challenges part of this section.
Figure 30: Two technologies behind edge intelligence.
IX-E2 INTELLIGENT EDGE
@@
  • Introduction: Intelligent edge, commonly known as edge intelligence (EI), refers to a combination of edge computing and AI [485], as shown in Figure 30. EI benefits from the synergy of both technologies. Edge computing drives computational capabilities, tasks, and applications from the core to the network edge [485]. The network edge consists of different types of edge nodes, typical ones being edge servers and end-devices (e.g., mobile and IoT devices) [485]. The edge computing paradigm shift, which moves computation resources closer to the edge information sources, naturally leads to reduced latency and communication overhead.
@@
  • Past and Present: While EI is a new concept, its components, AI and edge computing, have been studied separately in the past. AI research started already in the 1950s, with ups and downs on the road, until it reached a level of constant progress through the introduction of DL in the 2000s. Due to advances in other key technologies, AI boomed in the 2010s, reaching also mobile networks. AI was introduced for 5G in 3GPP Release 18 [16]. The history of AI was discussed in more detail in Section VIII. Although it took decades for AI to mature, edge computing has evolved rapidly. Specifically, edge computing became a popular topic in the 2010s, attracting wide interest in academia and industry. Consequently, edge computing has been supported by 5G since Release 15 [1], with constant enhancements in Release 16 [12], 17 [15], and 18 [16]. 3GPP is continuing its work on edge computing toward EI. The term EI was first introduced in the white paper of the Gartner Hype Cycle in 2018 [485]. Since then, EI has been one of the key 6G topics in academia and industry, spurring a broad range of publications, projects, collaborations, and development efforts.
@@
  • Opportunities and Challenges: Combining edge computing and AI/ML is a natural step in the evolution of mobile networks toward pervasive network intelligence. Edge computing offers a computation- and data-intensive platform, with the beneficial features of low latency, reduced communication overhead, high energy efficiency, protected privacy, and context awareness. AI has the ability to free the potential offered by edge computing by optimizing the network edge (e.g., performance, efficiency, adaptability, automation, and security/privacy) and enabling a vast range of disruptive applications. Due to its versatile benefits, EI is expected to play a major role in the 6G network intelligence. To realize large-scale EI in practice, fundamental challenges must be addressed and resolved. At a high level, the main challenges are related to the computation infrastructure, AI/ML models/algorithms, data availability, network management, real-time processing, network security, and privacy protection.
@@
    First, large-scale EI requires a ubiquitous computation infrastructure, with powerful computation and storage capabilities at the edge servers, which is a highly non-trivial trade-off between cost and performance. Given the stringent performance requirements of 6G, particularly in time-sensitive and mission-critical applications, real-time processing in AI/ML-driven network management is crucial to ensure fast and adaptable decision-making and network operations. This emphasizes the need for powerful computing, fast AI/ML algorithms, and efficient data acquisition mechanisms at the network edge. For each EI purpose/task, a dedicated AI/ML approach must be developed. Since EI will change the threat landscape, novel security approaches are required. Due to massive amounts of generated data, EI will be vulnerable to privacy threats, calling for advanced privacy protection techniques. For the success of EI, all the aforementioned issues must be properly addressed. Comprehensive discussions of open challenges and future directions can be found in [485, 486, 487, 488, 200].
@@
Figure 32: Three evolution phases toward AI-native air interface [500].
IX-E3 INTELLIGENT AIR INTERFACE
@@
    TABLE XI: Summary of beyond-communication technologies for 6G
@@
  • Literature and Future Directions: In the literature, the main research focus of AI/ML-enhanced air interface has been on DL-aided PHY layer designs, ranging from enhancing/replacing single or multiple processing blocks (such as channel coding, symbol mapping, channel estimation, equalization, detection, decoding, and symbol demapping) to replacing the entire PHY layer with an end-to-end learning process [500]. Furthermore, AI/ML-assisted MAC layer design has been widely examined, often related to resource allocation problems with DRL approaches [501]. Due to the immaturity of this field of research, major efforts are needed in all of these directions in the future, including AI/ML assistance from smaller to larger entities.
@@
    The AI/ML-assisted air interface has been reviewed in the recent literature [503, 504, 505, 506, 507, 508, 502, 500, 201, 509, 204, 510, 511]. In general, these surveys covered the main literature, recent advances, and open challenges. Most studies considered PHY layer design using DL [503, 504, 505, 506, 507, 508, 201, 204, 510, 511]. These DL works can be further divided into the 6G-specific [201, 204, 510], B5G/5G-specific [506, 508], and generic surveys [503, 504, 505, 507, 511]. Additionally, federated learning was reviewed for the PHY layer in [509]. The design of the MAC layer functions was discussed in [505, 502, 500]. The work [500] provided a vision toward an AI/ML-native 6G air interface, ideally enabling optimized communication schemes for any hardware, wireless environment, and application. The paper described the main transition phases on the road to achieving AI/ML-nativeness and discussed the needed PHY/MAC layer learning procedures. In [502], the authors introduced a framework for an AI/ML-enabled 6G air interface, covering PHY/MAC layer designs and interactions with the higher layers. Major research efforts are required in the near future to realize the vision of an intelligent 6G air interface. The corresponding challenges were discussed earlier. In particular, the 6G-specific characteristics and requirements must be properly addressed in the design of AI/ML-enhanced air interface.
IX-F BEYOND-COMMUNICATION TECHNOLOGIES FOR 6G
@@
The integration of beyond-communication technologies will be a major paradigm shift in the evolution of mobile networks, significantly extending the way how mobile networks can be exploited. The main beyond-communication technologies include computation, sensing, and energy, as shown in Figure 33. Currently, 5G networks have adopted edge computing capabilities and enhanced positioning in its palette of technologies. This paves the way for the broader usage of beyond-communication technologies in the 6G era. Beyond-communication technologies are expected to enable a broad range of novel capabilities and applications for 6G. In the following, we discuss integrated communication, computation, and caching (i3C), ISAC, and WET. These technologies are summarized in Table XI.
@@
IX-F1 INTEGRATED COMMUNICATION, COMPUTATION, AND CACHING
@@
  • Introduction: i3C refers to the joint design of 3C technologies at the edge of the network [512], as shown in Figure 34. Through joint design, 3C technologies can be used to benefit each other, enhancing the performance and extending the capabilities of mobile networks. Consequently, i3C enables novel services and applications. i3C is considered a major paradigm shift in the design of mobile networks [513].
@@
    Figure 34: Integrated 3C exploits the synergy between communication, computing, and caching at the network edge.
@@
  • Literature and Future Directions: In the literature, i3C has been reviewed from many different perspectives in a handful of survey papers [512, 515, 516, 517, 513, 514]. In [512], a survey was provided on mobile edge networks from the perspective of the 3C convergence. Computing and caching at the edge were first reviewed individually, and then the synergy between the 3C technologies was discussed. The work [515] discussed FL-assisted MEC, communication, and caching to provide more intelligence to the network edge. In [516], 3C resource sharing was explored for D2D IoT scenarios. The authors in [517] presented a review of the 3C convergence in IoT. In [513], a comprehensive survey was provided on the integration of 3C and control (i4C) for beyond 5G networks. The study [514] gave a thorough discussion on i3C in terms of cloud-edge cooperation, resource management, and intelligence.
@@
IX-F2 INTEGRATED SENSING AND COMMUNICATION
@@
    Due to the different nature of sensing and communication, cellular ISAC has certain shortcomings in practice [520]. First, ISAC requires a full-duplex operation or equivalent. Second, the sensing distance may be limited due to the limited transmission power of the cellular BS. Third, performance trade-offs exist between sensing and communication due to their conflicting targets and requirements. Typical trade-offs are defined as the PHY layer, spatial degrees of freedom, and cross-layer trade-offs, as discussed in [518].
@@
  • Literature and Future Directions: Since the late 2010s, ISAC has been studied from a vast range of aspects, such as information theoretical limits, performance trade-offs, communication-assisted sensing, sensing-assisted communication (beamforming, resource allocation, PHY layer security), signal processing (waveform optimization, MIMO design, receiver processing), network management (resource allocation, higher layer designs), integration with other emerging technologies (AI/ML, THz, RISs, V2X, UAVs, satellites), and potential applications (smart home/factory/city/healthcare, digital twins, XR, vehicular systems, remote sensing, environmental monitoring, etc.).
    The aforementioned topics and the related literature have been thoroughly reviewed in numerous surveys [520, 521, 522, 525, 519, 245, 524, 518, 526, 527, 528, 529, 530, 531, 251, 532, 533, 534, 535, 536, 537, 523, 255, 257, 538, 539, 540, 541, 542, 543, 544, 545]. At a high level, these survey papers can be classified into three categories: generic, specific, and 6G-oriented. In [520, 519, 524, 526, 535, 538], the generic surveys discussed ISAC from a wide perspective, covering topics such as fundamentals, state-of-the-art designs, recent advancements, potential applications, open problems, and future research guidelines. The specific surveys reviewed narrower aspects of ISAC, including fundamental limits [524], channel modeling [539], waveform design [527], ISAC signals [536], THz ISAC [245, 251, 255, 257], ISAC for IoT [525], RIS-assisted ISAC [531, 532, 533], holographic ISAC [530], UAV-assisted ISAC [534], ISAC for vehicular communication networks [528], and AI/ML-assisted ISAC [529, 537]. The 6G-oriented surveys provided comprehensive explorations of ISAC in the context of 6G [521, 522, 518, 523, 540, 541, 542, 543, 544, 545].
@@
IX-F3 WIRELESS ENERGY TRANSFER